Raja drives east from olds to trochu, a distance of 59 km. He then drives west to sundre, a distance of 97 km. What distance did raja drive? What was his total displacement?
1
Expert's answer
2016-05-06T13:41:02-0400
Distance is a scalar quantity, that is, l=59+97=156 (km),
Displacement is a vector quantity, that is, S=97-59=38(km), direction to the west.